graph the linear equation 8y-2x=-4 by finding and ploting its intercepts
Answer by rfer(16322)

8y-2x=-4
y int is when x=0
x int is when y=0
8y-2(0)=-4
y=-4/8
y=-1/2
y int= (0,-1/2)
----------------
8(0)-2x=-4
-2x=-4
x=2
x int= (2,0)